OpenStudy (anonymous):

Solve the equation involving rational expressions: x + 18/x = 11

OpenStudy (anonymous):

\[\frac{x+18}{x}=11\]??

OpenStudy (anonymous):

or its\[x+\frac{18}{x}=11\]

OpenStudy (anonymous):

Rayne are u there?

OpenStudy (anonymous):

I am here, sorry, work computers are awfully slow. So do I multiply x/x on both sides then?

OpenStudy (anonymous):

It is 18/x not the whole thing.

OpenStudy (anonymous):

\[x + \frac{ 18 }{ x } = 11\]

OpenStudy (anonymous):

so multiply by \[x^2-11x+18=0\]this is a quadratic solve by factoring or quadratic formula

OpenStudy (anonymous):

Mukushia, how did you come up with that?

OpenStudy (anonymous):

multiply both sides of equatoin by x \[x\times(x + \frac{ 18 }{ x }) = 11\times x\]\[x^2+x\times \frac{18}{x}=11x\]\[x^2+18=11x\]\[x^2-11x+18=0\]

OpenStudy (anonymous):

x^2+9x? That is what I got.

mathslover (mathslover):

Factorize that

OpenStudy (anonymous):

I got (x+3)^2, and then went onto find the x numbers, it was not the answer that was in the back of the book..

mathslover (mathslover):

\[\large{x^2-11x+18=0}\] \[\large{x^2-9x-2x+18=0}\] \[\large{x(x-9)-2(x-9)=0}\] \[\large{(x-2)(x-9)=0}\]

mathslover (mathslover):

that is either x = 2 or x = 9

OpenStudy (anonymous):

I see.. thank you.

mathslover (mathslover):

you're welcome